# Vikhr: The Family of Open-Source Instruction-Tuned Large Language Models for Russian

### v2

@@ 4.1 Experimental Setup
The evaluation was performed on MMLU Hendrycks et al. (2021), Ru-MMLU [^5] https://github.com/NLP-Core-Team/mmlu_ru, CheGeKa, Russian SuperGLUE Shavrina et al. (2020), and MERA Fenogenova et al. (2024). MMLU (En-MMLU) evaluates LLMs across 57 subjects with multiple-choice questions, assessing a model’s broad knowledge and reasoning abilities. We use this benchmark to verify that the model retains bi-lingual capabilities. In the results, we report the accuracy@1 score. RuMMLU is a translation of MMLU with GPT-3.5 to Russian. Just as for MMLU, we report the accuracy@1 score. CheGeKa is based on questions from the game “What? Where? When?”. This benchmark contains challenging open-ended questions, requiring logical reasoning and world knowledge. It includes 29,376 training and 416 test instances. The reported evaluation metric is the F1 score. Russian SuperGLUE is a benchmark similar to well-known English SuperGLUE Wang et al. (2019). It tests LLMs on various natural language understanding tasks like reading comprehension and textual entailment. The metric reported in the results is accuracy@1. The MERA benchmark encompasses 21 evaluation tasks for generative LLMs in 11 skill domains. Note that among other tasks MERA also includes CheGeKa, RuMMLU, and one of the subtasks of SuperGLUE (RWSD). The reported evaluation metric is the total score, which is the average of scores across all non-diagnostic tasks.
@@
We compare Vikhr to six open-source and two proprietary closed-source competitors of the similar size. Open-source models: aya101 – a massively multilingual LLM from CohereAI that follows instructions in 101 languages [^6] https://huggingface.co/CohereForAI/aya-101, it shows state-of-the-art results among massively multilingual LLMs; Mistral-7B-0.2-instruct – an English-oriented LLM that was used as the base model for Vikhr; rccmsu/ruadapt_mistral_saiga_7b_v0.1 – a Russian-oriented LLM that was constructed from the Mistral model using similar adaptations of the tokenizer, token embeddings, and the LM head Tikhomirov and Chernyshev (2023); saiga-mistral-7b-lora and saiga-llama3-8b – two versions of the Saiga models based on English-oriented LLMs and obtained by fine-tuning LoRA adapters on the Saiga instruction dataset [^7] https://huggingface.co/collections/IlyaGusev. Closed-source proprietary models for Russian: MTS AI Chat [^8] https://huggingface.co/MTSAIR/multi_verse_model and GigaChat-7b. The access to GigaChat weights is closed, so the reported results are taken from the leaderboards [^9] https://mera.a-ai.ru/ru/submits/10257. The results of MTS AI Chat are also taken from the leaderboard [^10] https://mera.a-ai.ru/ru/submits/10290.
4.2 Results
@@
We note that the original Mistral-7B-0.2-instruct, despite being an English-oriented model, demonstrates competitive performance in 3 out of 4 Russian benchmarks. This indicates demonstrates that such models could be viable alternatives. The only dataset, where its performance is very low is CheGeKa, which is related to open-ended question-answering. This may be due to the lack of culture-specific knowledge, as the English-oriented model has not seen much Russian texts. Note that the MTS AI Chat also shows very low results on CheGeKa, which might also indicate the lack of culture-specific knowledge.
The proprietary model GigaChat substantially outperforms Vikhr on CheGeKa and notably on Russian SuperGLUE. We assume this is due to the use of much larger Russian datasets for pre-training. However, surprisingly, it falls behind Vikhr on Ru-MMLU. On all benchmarks, Vikhr outperforms the the proprietary competitor from MTS AI.
@@ 5 Conclusion
We have presented Vikhr – a new state-of-the-art open-source instruction-following LLM oriented on the Russian language. To create Vikhr, we developed a comprehensive pipeline for adapting English-oriented LLMs to Russian. The pipeline includes the adaptation of the tokenizer vocabulary, continued pre-training of the entire model, and instruction tuning. We have also constructed a new dataset for instruction tuning by expanding the Saiga dataset with automatically translated and cleaned English instruction datasets. Our extensive work enabled Vikhr to outperform the known baselines, while maintaining computational efficiency.
We hope that the published models will foster the research on LLMs and enhance the diversity of languages incorporated into research agendas.

### v4

@@ 4.1 Experimental Setup
The evaluation was performed on MMLU (En-MMLU) Hendrycks et al. (2021), Ru-MMLU [^6] https://huggingface.co/datasets/NLPCoreTeam/mmlu_ru, Ru-MMLU-pro [^7] https://huggingface.co/datasets/Vikhrmodels/mmlupro-ru, ruXNLI Conneau et al. (2018), CheGeKa Mikhalkova and Khlyupin (2022), Russian SuperGLUE Shavrina et al. (2020), and MERA Fenogenova et al. (2024). The MMLU (En-MMLU) benchmark assesses LLMs across 57 subjects through multiple-choice questions, measuring a model’s general knowledge and reasoning abilities. We utilize this benchmark to verify that the model retains its bilingual proficiency. For this dataset, we report the accuracy@1 score. We constructed Ru-MMLU and Ru-MMLU-pro by automatically translating original English MMLU and MMLU-pro to Russian. Translation was done by GPT-3.5 and GPT-4 respectively. Just as for MMLU, for these datasets, we report the accuracy@1 score. CheGeKa is based on the questions from the Russian version of “Jeopardy” (the “Own game”). This benchmark contains challenging open-ended questions, requiring logical reasoning and world knowledge. It includes 29,376 training and 416 test instances. The reported evaluation metric is the F1 score. Russian SuperGLUE is a benchmark similar to well-known English SuperGLUE Wang et al. (2019). It tests LLMs on various natural language understanding tasks like reading comprehension and textual entailment. The metric reported in the results is accuracy@1. The MERA benchmark encompasses 21 evaluation tasks for generative LLMs in 11 skill domains. Note that among other tasks, MERA also includes CheGeKa, RuMMLU, and one of the subtasks of SuperGLUE (RWSD). The reported evaluation metric is the total score, which is the average of scores across all non-diagnostic tasks.
@@
We compare Vikhr to six open-source and two proprietary closed-source competitors of the similar size. Open-source models: aya101 Üstün et al. (2024) – a massively multilingual LLM from CohereAI that follows instructions in 101 languages [^8] https://huggingface.co/CohereForAI/aya-101, it shows state-of-the-art results among massively multilingual LLMs; Mistral-7B-0.2-instruct Jiang et al. (2023) – an English-oriented LLM that was used as the base model for Vikhr; rccmsu/ruadapt_mistral_saiga_7b_v0.1 Tikhomirov and Chernyshev (2023) – a Russian-oriented LLM that was constructed from the Mistral model using similar adaptations of the tokenizer, token embeddings, and the LM head; saiga-mistral-7b-lora and saiga-llama3-8b Gusev (2023) – two versions of the Saiga models based on English-oriented LLMs and obtained by fine-tuning LoRA adapters on the Saiga instruction dataset [^9] https://huggingface.co/collections/IlyaGusev. Closed-source proprietary models for Russian: MTS AI Chat [^10] https://huggingface.co/MTSAIR/multi_verse_model and GigaChat-7b. The access to GigaChat weights is closed, so the reported results are taken from the leaderboards [^11] https://mera.a-ai.ru/ru/submits/10257. The results of MTS AI Chat are also taken from the leaderboard [^12] https://mera.a-ai.ru/ru/submits/10290.
4.2 Results in NLU Tasks
@@
We note that the original Mistral-7B-0.2-instruct, despite being an English-oriented model, demonstrates competitive performance in 3 out of 4 Russian benchmarks. This demonstrates that such models could be viable alternatives at least for NLU tasks. The only dataset, where its performance is very low is CheGeKa, which is related to open-ended question-answering. This may be due to the lack of culture-specific knowledge, as the English-oriented model has not seen much Russian texts. Note that the MTS AI Chat also shows very low results on CheGeKa, which might also indicate the lack of culture-specific knowledge.
The proprietary model GigaChat substantially outperforms Vikhr on CheGeKa and notably on Russian SuperGLUE. We assume this is due to the use of much larger Russian datasets for pre-training. However, surprisingly, it falls behind Vikhr on Ru-MMLU. On all benchmarks, Vikhr outperforms the proprietary competitor from MTS AI.
[…]
33k 10.2 0.42
40k 14.4 0.46
60k 16.7 0.43
80k 20.4 0.41
Table 8: Performance of the intermediate Vikhr models with different vocabulary sizes after continued pre-training. The base model is LLaMa-2 7b. The vocabulary size selected for the final Vikhr model is underlined.
[…]
No filt. (17b tokens) 8.4 0.37 10.2
With filt. (11b tokens) (Ours) 7.2 0.46 11.1
Table 9: Performance metrics with and without filtration of the corpus for the continued pre-training. Perplexity is computed on the instruction dataset.
[…]
No regul. 8.1 0.34
With regul. (Ours) 7.2 0.45
Table 10: Performance with and without KL loss regularization during continued pre-training.
Instruction Set Ru-MMLU-pro
Saiga SFT 0.21
Translated Nectar 0.20
Ours 0.27
Table 11: Performance of Vikhr models fine-tuned on various instruction sets. The base model is Llama-3 8b.
4.3 Quality of Generated Text
Previous results demonstrate the performance of our model in NLU tasks. However, the performance in NLU does not always reflect the performance in text generation, as the former is based on classification and ranking capabilities of models.
To check the performance of the model in text generation, we performed qualitative analysis of LLM outputs. Table 7 compares several responses of Vikhr with outputs of the base model (Mistral-7B-Instruct-v0.2). As we can see from the presented examples, when Mistral generates Russian text, it often injects English words. Moreover, sometimes generated words consist of an English and a Russian token. From the second example, we see that Mistral also has issues with grammatical coherence. In the third example, LLM suggests to “clean a fish from fur”, which illustrates the lack of understanding of word meanings in the Russian language. We also note that Mistral tends to answer in English even when the input prompt is in Russian. These issues appear very often and make the base model useless for generation of Russian texts in practical scenarios. On the considered examples, Vikhr does not demonstrate any of these problems. Texts generated by Vikhr are grammatically coherent and correct.
4.4 Ablation Studies
We conducted several ablation studies to demonstrate effects of various features of our model translation pipeline that allowed Vikhr to achieve high performance.
Selection of the vocabulary size.
Table 8 presents the performance of intermediate Vikhr models with different vocabulary sizes after the stage of continued pre-training. We measure perplexity of the LLM on the instruction set and the performance on the ruXNLI task Conneau et al. (2018). As we can see, perplexity increases with the vocabulary size, indicating some degradation. However, the performance on the ruXNLI dataset is not monotonic. While the results for the largest vocabulary size are lower than those for the smallest, we observe a performance peak at a vocabulary size of 40k tokens. We selected this size as it offers improvements in the final task with only a slight increase in perplexity.
Effect of filtration of the corpus for continued pre-training
is illustrated in Table 9. As we can see, despite reducing the size of the data, performing continued pre-training on the filtered corpus results in a model with lower perplexity and substantially better scores in both considered end tasks: Ru-MMLU-pro and ruXNLI. This again highlights the importance of data quality for constructing good LLMs.
Effect of loss regularization in continued pre-training
is illustrated in Table 10. The results show that the KL regularization introduced in our pipeline slightly reduces perplexity and substantially increases the model performance in the ruXNLI task. This shows that continued pre-training on its own might deteriorate LLM reasoning capabilities, and proper regularization helps preventing the catastrophic forgetting.
Effect of fine-tuning on various instruction sets
is illustrated in Table 11. The results demonstrate that fine-tuning on our instruction set gives a big boost in performance on NLU tasks compared to Saiga and translated Nectar.
@@ 5 Conclusion
We have presented Vikhr – a new state-of-the-art open-source instruction-following bilingual LLM oriented on the Russian language. To create Vikhr, we developed a comprehensive pipeline for adapting English-oriented LLMs to other languages. The pipeline includes the adaptation of the tokenizer vocabulary, continued pre-training of the entire model, and instruction tuning. We have also constructed a new dataset for instruction tuning by expanding the Saiga dataset with automatically translated and cleaned English instruction datasets. Our extensive work enabled Vikhr to outperform the known baselines while maintaining computational efficiency.
We hope that the developed cross-lingual adaption pipeline and the published models will foster the research on LLMs and enhance the diversity of languages incorporated into research agendas.
In the future work, we plan to release in the open-source new better versions of Vikhr. At the moment, our best publicly available model is Vikhr-Nemo-12B-Instruct [^13] https://huggingface.co/Vikhrmodels/Vikhr-Nemo-12B-Instruct-R-21-09-24 based on Mistral NeMo. We also plan to perform cross-lingual adaption of LLMs to low-resource languages such as Belarusian, Serbian, and Kazakh.
@@
Ethical Considerations
The development and deployment of Vikhr raise several ethical considerations that must be addressed to ensure its responsible use:
  • Bias and Fairness: For developing Vikhr, we use publicly available data. Despite efforts to train Vikhr on diverse datasets, there is a risk of inherent biases in the data which may be reflected in the model’s outputs. Continued monitoring and evaluation are required to mitigate any biases, ensuring fair and unbiased performance.
  • Misinformation: As with any LLM, Vikhr has the potential to generate misleading or incorrect information. It is crucial to establish guidelines and mechanisms for users to verify the information provided by the model, promoting critical assessment and cross-referencing with reliable sources.
  • Misuse: Vikhr can be used for malicious purposes, such as generating harmful content, spam, or deepfakes. Implementing usage restrictions and monitoring mechanisms to detect and prevent misuse is critical to safeguard against these risks.
Acknowledgements
We thank the reviewers for their insightful comments, which have significantly improved this paper. We would like to express our sincere gratitude to Nikolay Kompanets for his invaluable contributions to the fine-tuning of the small models and the verification of hypotheses.